MARC (marc_d):

just to double check my answer @ganeshie8

MARC (marc_d):

\[If~\log_ax=1+\log_a(7x-10a)\]express x in terms of a.

MARC (marc_d):

i got x=3a/2 and x=-10a^2/(1-7a)

MARC (marc_d):

but i think the second answer is wrong

zepdrix (zepdrix):

Hmm I understand where the second answer came from... but where is this 3a/2 coming from?

MARC (marc_d):

actually,i hv two workings..i will show the working here..

MARC (marc_d):

\[\log_ax=\log_aa+\log_a(7x-10a)\]\[\log_ax-\log_a(7x-10a)=\log_aa\]\[\log_a(x-7x+10a)=\log_aa\]\[-6x=-9a\]\[x=\frac{ 3 }{ 2 }a\]

ganeshie8 (ganeshie8):

log(m) - log(n) is not same as log(m-n)

MARC (marc_d):

\[\log_ax-\log_a(7x-10a)=\log_aa\]\[\log_a \frac{ x }{ (7x-10a) }=\log_aa\]\[\frac{ x }{ 7x-10a }=a\]

MARC (marc_d):

\[x=-\frac{ 10a^2 }{ 1-7a }\]

MARC (marc_d):

This is my second working..

MARC (marc_d):

oh i c.. xD Thank you @ganeshie8 @zepdrix

ganeshie8 (ganeshie8):

yw :)